Please answer d,e,f and g, thank you!
roblem 1. Let (U common p.d.f. i 1 be a sequence of ii.d. discrete random variables with f(k) for k = 1, 2, 3 and for n 21 let Sn = Σ,u. (a) Find the probability that S2 is even. (b) Find the probability that Sn is even given that S,-1 is even. (e) Find the probability that Sn is even given that S-1 is odd. (d) Let pn P(Sn is even). Find pn in terms of pn-1. (e) Find the generating function for the pns. (f) Show that limn pn = 1 /2. (g) Suppose that the probability mass function of the U,s is not necessarily uniform but instead satisfies f(k) = Pk for k = 1 , 2, 3 where each Pk > 0 and Σ 1p.-1. What does the generating function for the pns look like now? In this more general setting does limn Pn 1/2?
Answer #1

(a) S2 will be even if both U1 and U2 are even or both U1 and U2 are odd.

P(S_2\texttt{ even}) = P(U_1\texttt{ even})\cdot P(U_2\texttt{ even}) + P(U_1\texttt{ odd})\cdot P(U_2\texttt{ odd}) \\ = \frac{1}{3}\cdot \frac{1}{3} + \frac{2}{3}\cdot \frac{2}{3} = \frac{5}{9}

(b) If Sn-1 is even, then we need Un to be even to make Sn even.

P(S_n\texttt{ even}|S_{n-1}\texttt{ even}) = P(U_n\texttt{ even}) = \frac{1}{3}

(c) If Sn-1 is odd, then we need Un to be odd to make Sn even.

P(S_n\texttt{ even}|S_{n-1}\texttt{ odd}) = P(U_n\texttt{ odd}) = \frac{2}{3}

(d) The probability of Sn being even will depend on the probability of Sn-1 being even.

\rho_n = \frac{1}{3}\cdot\rho_{n-1} + \frac{2}{3}\cdot(1-\rho_{n-1}) = \frac{2}{3}-\frac{1}{3}\cdot\rho_{n-1}=\frac{2-\rho_{n-1}}{3}

(e) We can calculate the few initial elements of this series. We get:

\rho_n = \left\{\frac{1}{3}, \frac{5}{9}, \frac{13}{27}, \frac{41}{81}, \cdots\right\}

We can observe that the denominators increase by a factor of 3 for every n. Also, the numerators are alternating between (3n-1)/2 for odd n and (3n+1)/2 for even n. From this, we get the generating functions as:

\rho_n = \left\{\begin{matrix} \frac{3^n-1}{2\cdot 3^n}=\frac{1}{2}-\frac{1}{2\cdot 3^n}, & n\texttt{ is odd}\\ \\ \frac{3^n+1}{2\cdot 3^n}=\frac{1}{2}+\frac{1}{2\cdot 3^n}, & n\texttt{ is even} \end{matrix}\right.

(f) We know that:

\lim_{n\to \infty}\frac{1}{3^n} = 0

Thus, both the series for \rho_n converge and we get:

\lim_{n\to\infty}\rho_n = \left\{\begin{matrix} \frac{1}{2}-\frac{1}{2}\lim_{n\to\infty}\frac{1}{3^n}=\frac{1}{2}, & n\texttt{ is odd}\\ \\ \frac{1}{2}+\frac{1}{2}\lim_{n\to\infty}\frac{1}{3^n}=\frac{1}{2}, & n\texttt{ is even} \end{matrix}\right.

So, both the series converge to the same limit and we get:

\lim_{n\to\infty}\rho_n = \frac{1}{2}

(g) In this case, the recursion formula will be:

\rho_n = p_2\cdot\rho_{n-1} + (p_1+p_3)\cdot(1-\rho_{n-1}) = p_2\cdot\rho_{n-1} + (1-p_2)\cdot(1-\rho_{n-1})\\ = (1-p_2)-(1-2\cdot p_2)\cdot\rho_{n-1}

We can see that the series converges by taking the difference between the successive terms:

\rho_n - \rho_{n-1} = \left[(1-p_2) - (1-2p_2)\rho_{n-1} \right ] -\left[(1-p_2) - (1-2p_2)\rho_{n-2} \right ] \\ = -(1-2p_2)(\rho_{n-1}-\rho_{n-2})

So, this series will be oscillating and will converge only if (1-2p2>0), i.e. if p2<1/2.